# Balanced Token Pruning: Accelerating Vision Language Models Beyond Local Optimization

### v1

@@ 6 Conclusion
In this work, we conduct initial studies to investigate and verify the limitations of existing image token pruning methods. We further analyze the impact of two pruning strategies on model performance from the perspective of the objective function, and formulate a local-global pruning optimization objective. To reduce information loss during pruning, we propose Balanced Token Pruning (BTP), a multi-stage pruning method. We first determine the pruning stages using a calibration set. In the early layers, we focus on a diversity-oriented objective to account for the influence of pruning on deeper layers, while in the later layers, we adopt an attention-based objective to better preserve local information. Experiments on several popular models and multiple visual understanding benchmarks demonstrate the effectiveness of our approach. We hope our method can inspire further research on visual token redundancy and the design of more efficient multimodal large models.
Appendix A Appendix
A.1 Key and Value of LVLMs
Following previous works on token quantization KIVI [24], we visualize the $K_{l}$ and $V_{l}$ of different LVLMs, the results are shown below:
(a) LLaVA key
(b) LLaVA value
Figure 7: Visualization of key and value of LLaVA-v1.5
(a) LLaVA-v1.6 key
(b) LLaVA-v1.6 value
Figure 8: Visualization of key and value of LLaVA-v1.6
(a) Qwen2.5-vl key
(b) Qwen2.5-vl value
Figure 9: Visualization of key and value of Qwen2.5-vl
A.2 Experiment Settings
For LLaVA-v1.5-7B, LLaVA-v1.5-13B, and LLaVA-v1.6-7B, we divide the pruning process into five stages based on the image token handling pipeline described in the Appendix. In each stage, except for the last one, we retain 50% of the tokens from the previous stage. In the final stage, all tokens are discarded to maximize inference speed. For Qwen2.5-VL, since its image token processing can be clearly divided into two stages, we retain 25% of the tokens in the fourth stage and 12.5% in the final stage to preserve model performance.
A.3 Visualization of token selection under different pruning strategies
Figure 10: Visualization of Image Token Selection Across Different Methods
A.4 Calculation of model gain
Since evaluation metrics vary across tasks and the difficulty levels differ significantly, it is not reasonable to present all task results directly in a unified format. For example, the original LLaVA-v1.5 model scores 1510 on the MME benchmark but only 62 on GQA. To address this, we define a model gain metric as:
$Gain=Normalize(\frac{Pruned_{score}}{Original_{score}}).$ (12)

### v2

@@ 6 Conclusion
In this work, we conduct initial studies to investigate and verify the limitations of existing image token pruning methods. We further analyze the impact of two pruning strategies on model performance from the perspective of the objective function, and formulate a local-global pruning optimization objective. To reduce information loss during pruning, we propose Balanced Token Pruning (BTP), a multi-stage pruning method. We first determine the pruning stages using a calibration set. In the early layers, we focus on a diversity-oriented objective to account for the influence of pruning on deeper layers, while in the later layers, we adopt an attention-based objective to better preserve local information. In future work, we will further investigate the lightweight deployment on real devices ruan2025edmambarethinkingefficientevent ; zheng2025reviewedgelargelanguage and explore its potential applications in multi-agent collaboration 10.1145/3594739.3612905 ; gao2025multimodalagenttuningbuilding .
Acknowledgments
This paper was supported by the Natural Science Foundation of China under Grant 62371269, Natural Science Foundation of China under 62272262, Shenzhen Low-Altitude Airspace Strategic Program Portfolio Z253061 and Guangdong Innovative, Entrepreneurial Research Team Program (2021ZT09L197) and Meituan Academy of Robotics Shenzhen. This paper was sponsored by Tsinghua University-Toyota Research Center.